OpenStudy (anonymous):

true or false? 7- square root of -100=17

OpenStudy (anonymous):

What number times itself is -100?

OpenStudy (anonymous):

im aginary number 10i

OpenStudy (anonymous):

false

OpenStudy (across):

yardbird84, remember that\[i=\sqrt{-1}.\]Therefore,\[\sqrt{-100}=\sqrt{(-1)(100)}=\sqrt{-1}\sqrt{100}=10i,\]as jimmyrep posted.